Determine if the function y = A sinkx cosωt is a solution to the wave equation.


1
Expert's answer
2021-10-24T18:23:23-0400

Given:

"y(x,t)=A\sin kx\cos \omega t"


The wave equation

"\frac{\partial^2y}{\partial t^2}-v^2\frac{\partial^2y}{\partial x^2}=0"


The speed of the wave "v=\omega/k"

We have


"\frac{\partial^2y}{\partial t^2}=-\omega^2A\sin kx\cos \omega t"

"\frac{\partial^2y}{\partial x^2}=-k^2A\sin kx\cos \omega t"

So

"-\omega^2A\sin kx\cos \omega t-v^2(-k^2A\sin kx\cos \omega t)="

"-k^2v^2A\sin kx\cos \omega t+k^2v^2A\sin kx\cos \omega t=0"